### v1

# Conformalized Polynomial Chaos Expansion for Uncertainty-aware Surrogate Modeling

Abstract
This work introduces a method to equip data-driven polynomial chaos expansion surrogate models with intervals that quantify the predictive uncertainty of the surrogate. To that end, we integrate jackknife-based conformal prediction into regression-based polynomial chaos expansions. The jackknife algorithm uses leave-one-out residuals to generate predictive intervals around the predictions of the polynomial chaos surrogate. The jackknife+ extension additionally requires leave-one-out model predictions. The key to efficient implementation is to leverage the linearity of the polynomial chaos regression model, so that leave-one-out residuals and, if necessary, leave-one-out model predictions can be computed with analytical, closed-form expressions, thus eliminating the need for repeated model re-training. In addition to the efficient computation of the predictive intervals, a significant advantage of this approach is its data efficiency, as it requires no hold-out dataset for prediction interval calibration, thus allowing the entire dataset to be used for model training. The conformalized polynomial chaos expansion method is validated on several benchmark models, where the impact of training data volume on the predictive intervals is additionally investigated.
Keywords: conformal prediction, jackknife, polynomial chaos expansion, regression, surrogate modeling, uncertainty quantification.
1.  Introduction
Originally developed as an efficient uncertainty quantification (UQ) method [1, 2, 3], the polynomial chaos expansion (PCE) is in recent years routinely employed as a machine learning (ML) regression model [4]. Within this context, regression-based PCEs have been used as data-driven surrogate models in various science and engineering applications [5, 6, 7, 8, 9]. Their use has also been extended to scientific ML tasks like manifold learning [10, 11, 12, 13], operator learning [14], and physics-informed ML [15, 16].
In contrast to the use of PCEs for UQ purposes, PCE-based data-driven predictive models are purely deterministic. That is, they yield point predictions that lack any type of uncertainty or confidence metrics, as opposed to Gaussian processes (GPs) [17] or Bayesian neural networks (NNs) [18], for example. Ensemble modeling and bootstrapping are common ways of providing such predictive uncertainty intervals to data-driven models, including PCEs [19, 20]. Other relevant approaches include combined PCE-GP models [21, 22] or stochastic PCE formulations [23]. However, these approaches induce either specific modeling limitations or significant computational overhead due to model retraining.
Considering predictive UQ for ML models, conformal prediction (CP) [24] is an established method for producing statistically valid prediction intervals given a point predictor, only under the assumption of data exchangeability. The resulting predictive models are commonly called “conformalized” [25, 26, 27, 28, 29]. CP is a distribution-free UQ method, meaning that it does not require any assumptions about the underlying data distribution or the structure of the data-driven model. Additionally, certain CP methods provide statistically rigorous, finite-sample guarantees regarding prediction interval coverage. On its downsides, CP does not yield full probability distributions like Bayesian methods do. Moreover, its coverage is marginal instead of conditional, hence possibly under- or overconfindent in different regions of the input space. Adaptive CP methods that address the latter issue are actively pursued [30, 31]. Data demand can also be an issue, in particular considering so-called inductive or split CP methods [32, 33], which require that part of the available data is excluded from training and reserved for prediction interval estimation. Remedies have been sought in so-called cross CP approaches leveraging cross-validation techniques [34, 35], including jackknife-based variants that utilize leave-one-out (LOO) residuals [33, 36]. Similar to bootstrapping and ensemble modeling, the downside of these methods is the need for repeated model retraining, which can result in undesired computational overhead.
This work proposes a novel combination of regression-based PCEs with CP based on the jackknife method [33] and its jackknife+ extension [36], for constructing predictive intervals that quantify the predictive uncertainty of data-driven PCEs. Combining CP with data-driven surrogate models has been explored before, e.g., for GPs [29] or various NNs architectures [37]. However, the use of regression-based PCEs results in an uncertainty-aware, data-driven surrogate modeling framework, which is additionally very computationally efficient. The key enabler is the linear form of PCE regression models, which allows to compute the LOO residuals required by the jackknife method analytically in closed form, without the need to re-train the PCE. The same is true for the LOO model predictions needed by the jackknife+ method. Due to this synergy of jackknife-based CP and PCE regression, the conformalized PCE method provides predictive intervals with minimal computational overhead. Additionally, the suggested approach belongs to the class of cross CP methods, thus allowing the full available dataset to be used for training. This is a crucial advantage in small-data learning regimes. To the authors’ knowledge, this is the first work to develop a conformalized PCE regression method.
The PCE method is implemented for total-degree PCEs utilizing ordinary least squares (OLS) for the computation of the PCE coefficients, and validated on four different benchmark models with one to ten inputs and scalar outputs. Extensions to other types of polynomial bases and coefficient estimation methods are conceptually straightforward, with the requirement for a fixed polynomial basis so that the closed-form formulas for the LOO residuals and model predictions remain valid (see section 2.3). Different training data availability conditions are also examined regarding their impact on prediction intervals and coverage. Prior works have shown the importance of training data volume for the accuracy and stability of regression-based PCEs [38, 39]. Here, we extend this investigation to assess how data availability impacts the quality of the predictive UQ provided by the conformalized PCE method.
The remaining of this paper is organized as follows. Section 2 presents the methodology behind the suggested method, first recalling the necessary prerequisites of PCEs and CP, and then presenting the conformalized PCE. The method is validated using several benchmark models in section 3. The paper concludes with section 4, which summarizes the findings of this work and discusses follow-up research opportunities.
2.  Methodology
We begin by presenting the PCE regression model in section 2.1, followed by a presentation of CP section 2.2. The latter emphasized the jackknife and jackknife+ CP methods. Section 2.3 shows how LOO residuals and model predictions of PCE regression models can be computed with analytical, closed-form expressions, as to enable the cost-effective estimation of predictive intervlas based on jackknife and jackknife+ CP. Last, section 2.4 discusses non-conformity score normalization for conformalized PCEs.
2.1 Regression-based polynomial chaos expansion
In the following we consider a computational model, denoted as $\mu(\mathbf{x})$, which receives an input vector $\mathbf{x}\in\mathbb{R}^{N}$ and computes a scalar output $y\in\mathbb{R}$, such that $y=\mu\left(\mathbf{x}\right)$. The scalar output is considered for simplicity only; vector-valued outputs can also be considered by an element-wise application of all operations described next. The form of the model can range from a simple analytical function to a complex numerical solver. The only assumption is that the model is deterministic, i.e., the same output $y$ is observed given the same input $\mathbf{x}$.
Next, we introduce a multivariate random variable (or random vector) $\mathbf{X}=\left(X_{1},\dots,X_{N}\right)$, where the individual random variables $X_{n}$, $n=1,\dots,N$, are assumed to be mutually independent. The random vector is defined on the probability space $\left(\Omega,\Sigma,P\right)$, where $\Omega$ is the sample space containing all measurable outcomes $\omega\in\Omega$, $\Sigma$ denotes the $\sigma$-algebra of events (grouped outcomes), and $P:\Sigma\rightarrow[0,1]$ is a probability measure. The latter is defined based on the assumed probability density function (PDF) $f_{\mathbf{X}}:\Xi\rightarrow\mathbb{R}_{\geq 0}$, where $\Xi\subset\mathbb{R}^{N}$ denotes the state-space of the random vector. Then, the random vector is a map, $\mathbf{X}:\Omega\rightarrow\Xi$, with random realizations $\mathbf{x}=\mathbf{X}\left(\omega\right)\in\Xi$.
Next, we assume that the random realizations $\mathbf{x}=\mathbf{X}\left(\omega\right)$ coincide with the inputs of the computational model $\mu(\mathbf{x})$ and that the Doob-Dynkin lemma is satisfied [40]. Then, by propagation of uncertainty, the model’s response is a random variable dependent on the input random vector, i.e., $Y=\mu\left(\mathbf{X}\right)$. Provided that its variance is finite, the random output can be represented in the form of a PCE, such that
$y=\mu\left(\mathbf{x}\right)=\sum_{\boldsymbol{\alpha}\in\mathbb{N}_{0}^{N}}c_{\boldsymbol{\alpha}}\Psi_{\boldsymbol{\alpha}}\left(\mathbf{x}\right),$ (1)
where $\Psi_{\boldsymbol{\alpha}}$ are multivariate polynomials, $c_{\boldsymbol{\alpha}}\in\mathbb{R}$ are the associated coefficients, and ${\boldsymbol{\alpha}}=\left(\alpha_{1},\dots,\alpha_{N}\right)$ are unique multi-indices defined by the polynomial’s partial degree in each input dimension. The polynomials are chosen such that they form an orthogonal basis with respect to the input PDF, i.e., they satisfy:
$\mathbb{E}\left[\Psi_{\boldsymbol{\alpha}}\Psi_{\boldsymbol{\beta}}\right]=\int_{\Xi}\Psi_{\boldsymbol{\alpha}}\left(\mathbf{x}\right)\Psi_{\boldsymbol{\beta}}\left(\mathbf{x}\right)f_{\mathbf{X}}\left(\mathbf{x}\right)\mathrm{d}\mathbf{x}=\mathbb{E}\left[\Psi_{\boldsymbol{\alpha}}^{2}\right]\delta_{\boldsymbol{\alpha}\boldsymbol{\beta}}.$ (2)
For simplicity, we assume an orthonormal basis, such that $\mathbb{E}\left[\Psi_{\boldsymbol{\alpha}}^{2}\right]=1$.
In practice, the infinite series (1) must be truncated to a finite number of $K$ terms based on a set of selected multi-indices. This truncation yields a PCE approximation of the form
$y=\mu\left(\mathbf{x}\right)\approx\widehat{\mu}\left(\mathbf{x}\right)=\sum_{k=1}^{K}c_{k}\Psi_{k}\left(\mathbf{x}\right)=\sum_{\boldsymbol{\alpha}\in\Lambda}c_{\boldsymbol{\alpha}}\Psi_{\boldsymbol{\alpha}}\left(\mathbf{x}\right),$ (3)
where $\Lambda\subset\mathbb{N}_{0}^{N}$ is a multi-index set with cardinality $\Lambda=K$. Each index $k=1,\dots,K$ corresponds to a unique multi-index $\boldsymbol{\alpha}\in\Lambda$. One common truncation method is to use a so-called total-degree basis, where $\Lambda=\left\{\boldsymbol{\alpha}:\left|\boldsymbol{\alpha}\right|_{1}\leq P\right\}$ and $P$ denotes the maximum polynomial degree. The size of the total-degree basis is equal to $K=\frac{\left(N+P\right)!}{N!P!}$, which can become intractable for high polynomial degrees or high-dimensional inputs, the latter being a manifestation of the so-called curse of dimensionality. To alleviate this problem, several methods have been suggested for constructing sparse bases [41], e.g., by means of dimension-adaptive algorithms [42].
Given a multi-index set $\Lambda$ and the corresponding polynomial basis, the remaining task is to estimate the PCE coefficients $\mathbf{c}=\left(c_{\boldsymbol{\alpha}}\right)_{\boldsymbol{\alpha}\in\Lambda}^{\top}\in\mathbb{R}^{K\times 1}$. We opt here for a regression-based approach [38, 39, 43]. Alternatives include include pseudo-spectral projection [44, 45] or interpolation [46, 47, 48]. For the regression-based coefficient estimation, we assume a set of input parameter realizations along with the corresponding model responses $\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$, commonly called the experimental design or, in an ML-context, the training data set. We then construct the so-called design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, such that $d_{mk}=\Psi_{k}\left(\mathbf{x}^{(m)}\right)$. Collecting the model responses in a vector $\mathbf{y}=\left(y^{(1)},\dots,y^{(M)}\right)^{\top}\in\mathbb{R}^{M\times 1}$, the coefficients can be computed by solving the least squares minimization problem
$\mathbf{c}=\underset{\hat{\mathbf{c}}\in\mathbb{R}^{K}}{\arg\min}\left\{\frac{1}{M}\lVert\mathbf{y}-\mathbf{D}\hat{\mathbf{c}}\rVert_{2}\right\},$ (4)
which admits the closed-form solution
$\mathbf{c}=\left(\mathbf{D}^{\top}\mathbf{D}\right)^{-1}\mathbf{D}^{\top}\mathbf{y}.$ (5)
For the regression problem (4) to be well posed, the experimental design must be at minimum equal to the polynomial basis. Larger experimental designs are typically needed for well determined problems. In fact, the size of the experimental design is crucial for the stability and accuracy of the resulting PCE [38, 39]. This limitation can be overcome with sparse PCE methods, where the problem (4) is complemented with a penalty term that induces sparsity in its solution. Common techniques include the least absolute shrinkage and selection operator (LASSO) and compressive sensing [41].
2.2 Conformal prediction
Following the notation introduced in section 2.1, we consider a regression model $\widehat{\mu}\left(\mathbf{x}\right)$ trained with the dataset $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. Then, given a test point $\left\{\mathbf{x}^{*},y^{*}\right\}\not\in\mathcal{D}_{\text{train}}$, the trained model produces a prediction $\widehat{y}^{*}=\widehat{\mu}(\mathbf{x}^{*})$ for the true value $y^{*}$. The goal of CP is to complement the model’s prediction with an interval which is likely to contain the true output $y^{*}$. More formally, given a significance level $s\in\left(0,1\right)$ corresponding to the target coverage level (1s)(1-s), CP constructs a prediction interval $\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)$, such that
$\mathbb{P}\left\{y^{*}\in\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)\right\}\geq 1-s.$ (6)
That is, the probability of the true output being included in the interval is at least equal to the target coverage level. Note that CP assumes that the training and test data are exchangeable, meaning that the joint data distribution is invariant to permutation.
[…]
$\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)\;=\;\left\{\,y:\;\alpha\left(\mathbf{x}^{*},y;\widehat{\mu}\right)\leq\widehat{q}_{1-s}\left\{\alpha_{j}\right\}\,\right\},$ (7)
such that $y^{*}\in\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)$ implies that $\alpha\!\left(\mathbf{x}^{*},y^{*};\widehat{\mu}\right)\leq\widehat{q}_{1-s}\!\left\{\alpha_{j}\right\}$ and vice versa. One common non-conformity score is the absolute residual, such that $\alpha_{j}=\left|r_{j}\right|=\left|y^{(j)}-\widehat{\mu}\left(\mathbf{x}^{(j)}\right)\right|$. This choice results in the construction of the symmetric predictive interval
$\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)=\left[\widehat{\mu}\left(\mathbf{x}^{*}\right)-\widehat{q}_{1-s}\left\{\left|r_{j}\right|\right\},\widehat{\mu}\left(\mathbf{x}^{*}\right)+\widehat{q}_{1-s}\left\{\left|r_{j}\right|\right\}\right].$ (8)
Note that this predictive interval does not generally adapt to possible data heteroskedasticity.
A naive approach to estimate the predictive interval (8) would be to use the residuals on the training data $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. However, training data residuals are expected to be smaller compared to residuals obtained from previously unseen test data points. Hence, the resulting predictive interval will most probably not provide the required coverage. Split or inductive CP approaches [32, 33] use a separate validation dataset $\mathcal{D}_{\text{val}}\neq\mathcal{D}_{\text{train}}$ to compute the non-conformity scores and construct the prediction interval (8). Split CP guarantees the required finite-sample coverage (6), however, this comes at the cost of larger data demand due to the need for a validation dataset. To avoid the need for additional data, cross CP [34, 35] utilizes $K$-fold cross-validation, such that the training dataset $\mathcal{D}_{\text{train}}$ is split into $K$ disjoint subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{K}$ of (approximately) equal size. Then, the regression models $\widehat{\mu}_{\sim k}\left(\mathbf{x}\right)$ are trained using the datasets $\mathcal{D}_{\sim k}=\mathcal{D}_{\text{train}}\setminus\mathcal{D}_{k}$, $k=1,\dots,K$, and the cross-validation residuals
$r_{m}^{\text{cv}}=y^{(m)}-\widehat{\mu}_{\sim k(m)}\left(\mathbf{x}^{(m)}\right),$ (9)
are used as non-conformity scores (in absolute value), where the index $k(m)$ denotes that $\left(\mathbf{x}^{(m)},y^{(m)}\right)\in\mathcal{D}_{k}$. Cross CP typically yields an empirical coverage close to the target level (1s)(1-s), but does not provide formal coverage guarantees. A major downside of cross CP is that $K$ different regression models must be trained, which can lead to an undesired computational cost.
A special case of cross-validation is the jackknife method [49]. Using the jackknife in the context of CP [33], the non-conformity scores are computed using the LOO residuals
$r_{m}^{\text{LOO}}=y^{(m)}-\widehat{\mu}_{\sim m}\left(\mathbf{x}^{(m)}\right),$ (10)
where the regression model $\widehat{\mu}_{\sim m}\left(\mathbf{x}\right)$ is trained using the dataset $\mathcal{D}_{\sim m}=\mathcal{D}_{\text{train}}\setminus\left\{\left(\mathbf{x}^{(m)},y^{(m)}\right)\right\}$. Using the absolute LOO residuals as non-conformity scores, the predictive interval (8) takes two equivalent forms, i.e.,
$\widehat{C}_{1-s}^{\text{jack}}\left(\mathbf{x}^{*}\right)$ $=\left[\widehat{\mu}\left(\mathbf{x}^{*}\right)-\widehat{q}_{1-s}\left\{\left|r_{m}^{\text{LOO}}\right|\right\},\widehat{\mu}\left(\mathbf{x}^{*}\right)+\widehat{q}_{1-s}\left\{\left|r_{m}^{\text{LOO}}\right|\right\}\right]$ (11a)
$=\left[\widehat{q}_{s}\left\{\widehat{\mu}\left(\mathbf{x}^{*}\right)-\left|r_{m}^{\text{LOO}}\right|\right\},\widehat{q}_{1-s}\left\{\widehat{\mu}\left(\mathbf{x}^{*}\right)+\left|r_{m}^{\text{LOO}}\right|\right\}\right].$ (11b)
Being a cross CP variant, jackknife CP also yields an empirical coverage close to the target level, also without formal coverage guarantees. Barber et al. [36] suggested an extension of jackknife CP, called jackknife+, where the predictive interval is constructed as
$C_{1-s}^{\text{jack+}}\left(\mathbf{x}^{*}\right)=\left[\widehat{q}_{s}\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)-\left|r_{m}^{\text{LOO}}\right|\right\},\widehat{q}_{1-s}\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)+\left|r_{m}^{\text{LOO}}\right|\right\}\right].$ (12)
In contrast to the jackknife interval (11) which is centered at the prediction $\widehat{\mu}\left(\mathbf{x}^{*}\right)$ of the fully trained model, the jackknife+ interval (12) is based on the ensemble of the LOO predictions $\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)$, $m=1,\dots,M$. Hence, the jackknife+ predictive interval is asymmetric. This slight modification results in a formal coverage guarantee of $\left(1-2s\right)$, while the empirical coverage remains close to the target level (1s)(1-s), similar to cross CP and jackknife CP.
It should be noted that, in practice, the jackknife and jackknife+ CP methods produce almost identical predictive intervals and both achieve an empirical coverage close to the target level (1s)(1-s) [36]. Moreover, both jackknife-based approaches entail training $M$ regression models, thus being the extreme case of cross CP in terms of computational cost. However, for linear regression models like the PCE, the LOO residuals $r_{m}^{\text{LOO}}$ and the LOO predictions $\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)$ can be obtained without model retraining, as shown in section 2.3.
2.3 Conformalized polynomial chaos expansion
We now assume that $\widehat{\mu}\left(\mathbf{x}\right)$ is a regression-based PCE in the form of (3) and trained upon a dataset $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. Then, the predictive intervals (11) (for jackknife) and (12) (for jackknife+) can be estimated without the need to retrain the PCE $M$ times.
Starting with jackknife method, only the LOO residuals defined in (10) are needed to estimate the prediction interval (11). Recalling from section 2.1 the design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, we define the so-called hat matrix $\mathbf{H}\in\mathbb{R}^{M\times M}$ as
$\mathbf{H}=\mathbf{D}\left(\mathbf{D}^{\top}\mathbf{D}\right)^{-1}\mathbf{D}^{\top}.$ (13)
The LOO residuals can then be computed analytically via the closed-form expression
$r_{m}^{\text{LOO}}=\frac{y^{(m)}-\widehat{\mu}\left(\mathbf{x}^{(m)}\right)}{1-h_{mm}},$ (14)
where $h_{mm}\in\mathbb{R}$ is the $m$-th diagonal term of $\mathbf{H}$. The derivation of the analytical expression (14) is shown in appendix A.
For the predictive interval (12) based on the jackknife+, the LOO predictions $\left\{\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)\right\}_{m=1}^{M}$ for the test input $\mathbf{x}^{*}$ are additionally needed. Let us denote with $\mathbf{d}_{m}^{\top}\in\mathbb{R}^{1\times K}$ the $m$-th row of the design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, which contains the evaluation of the PCE basis upon the training input $\mathbf{x}^{(m)}$, such that $\mathbf{d}_{m}^{\top}=\left(\Psi_{1}\left(\mathbf{x}^{(m)}\right),\dots,\Psi_{K}\left(\mathbf{x}^{(m)}\right)\right)$. Accordingly, let us denote with $\mathbf{d}_{*}^{\top}\in\mathbb{R}^{1\times K}$ the vector of basis evaluations at the test input $\mathbf{x}^{*}$, such that $\mathbf{d}_{*}^{\top}=\left(\Psi_{1}\left(\mathbf{x}^{*}\right),\dots,\Psi_{K}\left(\mathbf{x}^{*}\right)\right)$. Then, the LOO predictions can be analytically computed as
$\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)$ $=\widehat{\mu}\left(\mathbf{x}^{*}\right)-\mathbf{d}_{*}^{\top}\left(\mathbf{D}^{\top}\mathbf{D}\right)^{-1}\mathbf{d}_{m}\frac{y^{(m)}-\widehat{\mu}\left(\mathbf{x}^{(m)}\right)}{1-h_{mm}}$
$=\widehat{\mu}\left(\mathbf{x}^{*}\right)-\mathbf{d}_{*}^{\top}\left(\mathbf{D}^{\top}\mathbf{D}\right)^{-1}\mathbf{d}_{m}r_{m}^{\text{LOO}}.$ (15)
The derivation of the analytical expression (2.3) is shown in appendix B.
The analytical expressions (14) and (2.3) enable the efficient computation of jackknife and jackknife+ predictive intervals for PCE-based regression models, without the need for repeated model retraining. Note that these closed-form relations are valid for PCEs with a fixed polynomial basis. Should the basis be adaptively constructed, for example by means of sparsity-promoting algorithms [41, 42], explicit model refitting is required during the basis selection procedure. However, once the final basis is available, the analytical relations remain applicable.
2.4 Normalized non-conformity scores
In the previous section, the non-conformity scores were equal to the absolute LOO residuals, i.e., $\alpha_{m}=\left|r_{m}^{\text{LOO}}\right|$, $m=1,\dots,M$. Using the relative LOO error, defined as
$\epsilon^{\text{LOO}}_{\text{rel}}=\frac{\epsilon^{\text{LOO}}}{\mathrm{Var\left(Y\right)}}=\frac{\frac{1}{M}\sum_{m=1}^{M}\left(r_{m}^{\text{LOO}}\right)^{2}}{\mathrm{Var\left(Y\right)}},$ (16)
where $\mathrm{Var(Y)}$ denotes the variance of the output, normalized LOO-residual-based non-conformity scores can be defined as
$\alpha_{m}^{\text{norm}}=\frac{\left|r_{m}^{\text{LOO}}\right|}{\sqrt{\text{Var(Y)}}}.$ (17)
Then, the quantile of the CP interval is taken on these dimensionless scores and later rescaled as
$q_{1-s}^{\text{orig}}=q_{1-s}^{\text{norm}}\sqrt{\text{Var}(Y)}.$ (18)
The main advantage of using normalized non-conformity scores is that they adapt to heteroskedastic data, thus offering prediction intervals that vary in width according to local uncertainty. The main downside is that an output variance estimate is required, which may be challenging. For a PCE model, the output variance can be estimated cost-effectively by post-processing its coefficients, such that
$\mathrm{Var}(Y)\approx\sum_{\boldsymbol{\alpha}\in\Lambda\setminus\mathbf{0}}c_{\mathbf{\alpha}}^{2},$ (19)
assuming an orthonormal polynomial basis. Note that this estimate is affected by the accuracy of the PCE. Alternatively, the ouput variance can be estimated from the experimental design, assuming that there exist enough data points to allow for a sufficiently accurate estimate.
3.  Numerical experiments
The conformalized PCE is applied to benchmark models that are commonly used for the validation of surrogate modeling methods. A univariate model is used in section 3.1, while multivariate models with six, seven, and ten input dimensions are used in sections 3.2, 3.3, and 3.4, respectively. The multivariate models are taken from the “Virtual Library of Simulation Experiments: Test Functions and Datasets” by Surjanovic and Bingham [50]. Both the jackknife and jackknife+ methods are employed, with either normalized or non-normalized non-conformity scores. Additionally, we investigate the capability of the conformalized PCE method to construct predictive intervals that provide the target coverage, depending on training data availability. In all cases, the experimental design used to train the PCE consists of samples drawn randomly from the distribution of the input parameters. After model training and predictive interval estimation, the conformalized PCE is compared against the true function based on a separate test dataset. The same test dataset is employed to estimate the empirical coverage. Each experiment is repeated for $100$ random seeds to estimate coverage and predictive interval statistics. For all test-cases, the target coverage is fixed at $95\%$ and the test dataset consists of $10^{4}$ random samples. A discussion based on the numerical results is available in section 3.5.
3.1 Meromorphic function
We consider the one-dimensional meromorphic function
$\mu(x)=\frac{1}{a+bx},$ (20)
with $a=1$ and $b=0.5$. We assume that the input $x$ takes values as uniformly distributed in $\left[-1,1\right]$. The conformalized PCE is constructed with maximum polynomial degrees $P\in\left\{2,3\right\}$. Since the input is one-dimensional, the polynomial basis consists of $P+1$ terms. The experimental design has size $M=C\left(P+1\right)^{2}$, where $C\in\left\{2,3,5,10\right\}$ is an oversampling coefficient. Note that the quadratic scaling of the experimental design in relation to the polynomial basis is necessary to ensure a well conditioned least squares problem (4) in the one-dimensional input case [39].
Figure 1 compares the different configurations of the conformalized PCE against the true meromorphic function for a single random seed. Figure 2 shows the empirical coverage results, where the box plots summarize the distribution of the empirical coverage over the $100$ random seeds. Similarly, Figure 3 shows box plots regarding the corresponding predictive intervals. Note that, for this particular test-case, non-conformity score normalization has a negligible impact on coverage and predictive interval width, therefore only one set of results is presented.
(a) $P=2$, $C=2$.
(b) $P=2$, $C=3$.
(c) $P=2$, $C=5$.
(d) $P=2$, $C=10$.
(e) $P=3$, $C=2$.
(f) $P=3$, $C=3$.
(g) $P=3$, $C=5$.
(h) $P=3$, $C=10$.
Figure 1: Parity plots comparing ground truth values of the meromorphic function against conformalized PCE predictions for different combinations of polynomial degree $P$ and oversampling coefficient $C$. The results correspond to a single random seed. The results with and without non-conformity score normalization are identical.
(a) $P=2$.
(b) $P=3$.
Figure 2: Box plots of the empirical coverage provided by conformalized PCE surrogates of the meromorphic function, for different combinations of polynomial degree $P$ and oversampling coefficient $C$. The results with and without non-conformity score normalization are identical.
(a) $P=2$.
(b) $P=3$.
Figure 3: Box plots of the predictive interval widths provided by conformalized PCE surrogates of the meromorphic function, for different combinations of polynomial degree $P$ and oversampling coefficient $C$. The results with and without non-conformity score normalization are identical.
3.2 OTL circuit function
We consider the response of an output transformerless (OTL) push-pull circuit, given as
$\mu(\mathbf{x})=\frac{\left(\tfrac{12\,R_{b2}}{R_{b1}+R_{b2}}+0.74\right)\,\beta\,\left(R_{c2}+9\right)}{\beta\,(R_{c2}+9)+R_{f}},\;+\;\frac{11.35\,R_{f}}{\beta\,\left(R_{c2}+9\right)+R_{f}}\;+\;\frac{0.74\,R_{f}\,\beta\,(R_{c2}+9)}{\left(\beta\,\left(R_{c2}+9\right)+R_{f}\right)\,R_{c1}},$ (21)
where the the input vector $\mathbf{x}\in\mathbb{R}^{6}$ consists of the five resistances $R_{b1}$, $R_{b2}$, $R_{f}$, $R_{c1}$, $R_{c2}$, and the current gain $\beta$. The response corresponds to the midpoint voltage. The input parameters are assumed to be uniformly distributed within the value ranges given in Table 1.
Table 1: Input parameters of the OTL circuit function.
Parameter Units Range
[…]
$\beta$ AA $\left[50,300\right]$
Total-degree conformalized PCEs with maximum polynomial degrees $P\in\left\{1,2,3\right\}$ and oversampling coefficients $C\in\left\{2,3,5,10\right\}$ are employed. Following Migliorati et al. [39], the experimental design now scales linearly with the size of the PCE basis, such that $M=CK=C\#\Lambda$. The different conformalized PCE configurations are compared against the true OTL circuit function for a single random seed in Figure 4. This plot does not compare normalized versus non-normalized conformity scores, as the differences between the corresponding predictive intervals are not easily discernible. Figures 5 and 6 shows box plots regarding the distributions of the conformalized PCEs’ empirical coverages and predictive intervals over the 100 random seeds, also including the two different non-conformity scores.
(a) $P=1$, $C=2$.
(b) $P=1$, $C=3$.
(c) $P=1$, $C=5$.
(d) $P=1$, $C=10$.
(e) $P=2$, $C=2$.
(f) $P=2$, $C=3$.
(g) $P=2$, $C=5$.
(h) $P=2$, $C=10$.
(i) $P=3$, $C=2$.
(j) $P=3$, $C=3$.
(k) $P=3$, $C=5$.
(l) $P=3$, $C=10$.
Figure 4: Parity plots comparing ground truth values of the OTL circuit function against conformalized PCE predictions for different combinations of polynomial degree $P$ and oversampling coefficient $C$. The results correspond to a single random seed. The results with and without non-conformity score normalization are very similar, therefore, only one set of results in shown.
(a) $P=1$, $\alpha_{m}$.
(b) $P=2$, $\alpha_{m}$.
(c) $P=3$, $\alpha_{m}$.
(d) $P=1$, $\alpha_{m}^{\text{norm}}$.
(e) $P=2$, $\alpha_{m}^{\text{norm}}$.
(f) $P=3$, $\alpha_{m}^{\text{norm}}$.
Figure 5: Box plots of the empirical coverage provided by conformalized PCE surrogates of the OTL circuit function, for different combinations of polynomial degree $P$, oversampling coefficient $C$, and non-conformity score type.
(a) $P=1$, $\alpha_{m}$.
(b) $P=2$, $\alpha_{m}$.
(c) $P=3$, $\alpha_{m}$.
(d) $P=1$, $\alpha_{m}^{\text{norm}}$.
(e) $P=2$, $\alpha_{m}^{\text{norm}}$.
(f) $P=3$, $\alpha_{m}^{\text{norm}}$.
Figure 6: Box plots of the predictive interval widths provided by conformalized PCE surrogates of the OTL circuit function, for different combinations of polynomial degree $P$, oversampling coefficient $C$, and non-conformity score type.
3.3 Piston simulation function
We consider the circular motion of a piston within a cylinder, the cycle-time of which is given by the function
$\mu(\mathbf{x})=2\pi\sqrt{\frac{M}{k+S^{2}\frac{P_{0}V_{0}}{T_{0}}\frac{T_{0}}{V^{2}}}},$ (22)
$\text{where }V=\frac{S}{2k}\left(\sqrt{A^{2}+4k\frac{P_{0}V_{0}}{T_{0}}T_{a}}-A\right)\text{ and }A=P_{0}\,S+19.62\,M-\frac{k\,V_{0}}{S},$
where the input vector $\mathbf{x}\in\mathbb{R}^{7}$ consists of the parameters listed in Table 2 along with their value ranges. All parameters are assumed to be uniformly distributed within their ranges.
Table 2: Input parameters of the piston simulation function.
Parameter Description Units Range
$M$ piston weight kg $\left[30,60\right]$
$S$ piston surface area m2 $\left[0.005,0.02\right]$
$V_{0}$ initial gas volume m3 $\left[0.002,0.01\right]$
$k$ spring coefficient N/m $\left[1000,5000\right]$
$P_{0}$ atmospheric pressure N/m2 $\left[90000,110000\right]$
$T_{a}$ ambient temperature K $\left[290,296\right]$
$T_{0}$ filling gas temperature K $\left[340,360\right]$
Total-degree conformalized PCEs with maximum polynomial degrees $P\in\left\{2,3,4\right\}$ and oversampling coefficients $C\in\left\{2,3,5,10\right\}$ are employed. Similar to section 3.2, the experimental design scales linearly with the size of the PCE basis, i.e., $M=CK=C\#\Lambda$. Figure 7 compares the various configurations of the conformalized PCE against the true piston simulation function for a single random seed. Figures 8 and 9 show coverage and predictive interval statistics over the 100 random seeds.
(a) $P=2$, $C=2$.
(b) $P=2$, $C=3$.
(c) $P=2$, $C=5$.
(d) $P=2$, $C=10$.
(e) $P=3$, $C=2$.
(f) $P=3$, $C=3$.
(g) $P=3$, $C=5$.
(h) $P=3$, $C=10$.
(i) $P=4$, $C=2$.
(j) $P=4$, $C=3$.
(k) $P=4$, $C=5$.
(l) $P=4$, $C=10$.
Figure 7: Parity plots comparing ground truth values of the piston simulation function against conformalized PCE predictions for different combinations of polynomial degree $P$ and oversampling coefficient $C$. The results correspond to a single random seed. The results with and without non-conformity score normalization are very similar, therefore, only one set of results in shown.
(a) $P=2$, $\alpha_{m}$.
(b) $P=3$, $\alpha_{m}$.
(c) $P=4$, $\alpha_{m}$.
(d) $P=2$, $\alpha_{m}^{\text{norm}}$.
(e) $P=3$, $\alpha_{m}^{\text{norm}}$.
(f) $P=4$, $\alpha_{m}^{\text{norm}}$.
Figure 8: Box plots of the empirical coverage provided by conformalized PCE surrogates of the piston simulation function, for different combinations of polynomial degree $P$, oversampling coefficient $C$, and non-conformity score type.
(a) $P=2$, $\alpha_{m}$.
(b) $P=3$, $\alpha_{m}$.
(c) $P=4$, $\alpha_{m}$.
(d) $P=2$, $\alpha_{m}^{\text{norm}}$.
(e) $P=3$, $\alpha_{m}^{\text{norm}}$.
(f) $P=4$, $\alpha_{m}^{\text{norm}}$.
Figure 9: Box plots of the predictive interval widths provided by conformalized PCE surrogates of the piston simulation function, for different combinations of polynomial degree $P$, oversampling coefficient $C$, and non-conformity score type.
3.4 Wing weight function
We consider a function that models the weight of light aircraft wing, given as
$\mu(\mathbf{x})=0.036S_{\text{w}}^{0.758}W_{\text{fw}}^{0.0035}\left(\frac{A}{\cos^{2}(\Lambda)}\right)^{0.6}q^{0.006}\lambda^{0.04}\left(\frac{100t_{\text{c}}}{\cos(\Lambda)}\right)^{-0.3}(N_{\text{z}}\ W_{\text{dg}})^{0.49}+S_{\text{w}}W_{\text{p}},$ (23)
where the input vector $\mathbf{x}\in\mathbb{R}^{10}$ consists of the parameters listed in Table 3, along with their value ranges. All parameters are assumed to be uniformly distributed within their ranges.
Table 3: Input parameters of the wing weight function.
Parameter Description Units Range
$S_{\text{w}}$ wing area ft2 $\left[150,200\right]$
$W_{\text{fw}}$ wing fuel weight lb $\left[220,300\right]$
AA aspect ratio - $\left[6,10\right]$
$\Lambda$ quarter-chord sweep $(^{o})$ $\left[-10,10\right]$
$q$ dynamic pressure at cruise lb/ft2 $\left[16,45\right]$
$\lambda$ taper ratio - $\left[0.5,1\right]$
$t_{\text{c}}$ airfoil thickness to chord ratio - $\left[0.08,0.18\right]$
$N_{\text{z}}$ ultimate load factor - $\left[2.5,6\right]$
$W_{\text{dg}}$ flight design gross weight lb $\left[1700,2500\right]$
$W_{\text{p}}$ paint weight lb/ft2 $\left[0.025,0.08\right]$
Total-degree conformalized PCEs with maximum polynomial degrees $P\in\left\{1,2\right\}$ and oversampling coefficients $C\in\left\{2,3,5,10\right\}$ are employed. The experimental design scales linearly with the size of the PCE basis, i.e., $M=CK=C\#\Lambda$. Figure 10 compares the different conformalized PCE configurations against the true wing weight function for a single random seed. Figures 8 and 9 present coverage and predictive interval statistics over the 100 random seeds.
(a) $P=1$, $C=2$.
(b) $P=1$, $C=3$.
(c) $P=1$, $C=5$.
(d) $P=1$, $C=10$.
(e) $P=2$, $C=2$.
(f) $P=2$, $C=3$.
(g) $P=2$, $C=5$.
(h) $P=2$, $C=10$.
Figure 10: Parity plots comparing ground truth values of the wing weight function against conformalized PCE predictions for different combinations of polynomial degree $P$ and oversampling coefficient $C$. The results correspond to a single random seed. The results with and without non-conformity score normalization are very similar, therefore, only one set of results in shown.
(a) $P=1$, $\alpha_{m}$.
(b) $P=1$, $\alpha_{m}^{\text{norm}}$.
(c) $P=2$, $\alpha_{m}$.
(d) $P=2$, $\alpha_{m}^{\text{norm}}$.
Figure 11: Box plots of the empirical coverage provided by conformalized PCE surrogates of the wing weight function, for different combinations of polynomial degree $P$, oversampling coefficient $C$, and non-conformity score type.
(a) $P=1$, $\alpha_{m}$.
(b) $P=1$, $\alpha_{m}^{\text{norm}}$.
(c) $P=2$, $\alpha_{m}$.
(d) $P=2$, $\alpha_{m}^{\text{norm}}$.
Figure 12: Box plots of the predictive interval widths provided by conformalized PCE surrogates of the wing weight function, for different combinations of polynomial degree $P$, oversampling coefficient $C$, and non-conformity score type.
3.5 Discussion on the numerical results
The numerical results presented in sections 3.1-3.4 show a number of patterns and similarities. First, the conformalized PCE is in general capable of providing empirical coverages close to the target level, more commonly above it. For less accurate PCEs, e.g., ones corresponding to low polynomial degrees and/or small experimental designs, the coverage over the 100 random seeds varies significantly. As the accuracy of the PCE increases with larger experimental designs and higher polynomial degrees, the variability in coverage is reduced and the coverage moves closer to the target level.
The coverage provided by the conformalized PCE is naturally connected to the corresponding predictive intervals. For the less accurate PCEs, the required coverage is accomplished by larger predictive intervals, which also show a large variability over the 100 random seeds. As expected, the predictive intervals become smaller as the accuracy of the PCE increases. Importantly, even for low-degree PCEs, increasing the size of the experimental design leads not only to a reduced variability in the predictive interval, but also to a reduced interval width.
Comparing the jackknife and jackknife+ methods, their predictive intervals are in general very similar, especially for larger experimental designs and more accurate PCEs. The jackknife method is typically more conservative and results in comparatively larger intervals and to over-coverage, especially for smaller experimental designs. On the other hand, jackknife+ produces empirical coverages closer to the target level, in most cases. For both jackknife and jackknife+, it is clear that training data availability plays a significant role.
Last, comparing normalized and non-normalized non-conformity scores, no concrete conclusions can be made based on the available numerical results. There exist differences between the two, which are more common for PCEs trained with small experimental designs. However, these differences are generally small and not consistent. Overall, it can be said that both non-conformity scores perform similarly.
4.  Summary, conclusions, and outlook
This work presented a new method that combines regression-based PCEs with jackknife-based CP for constructing uncertainty-aware, data-driven surrogate models. Using this method, the data-driven PCE is complemented with a predictive interval that quantifies the uncertainty in its predictions. Crucially, this predictive interval is estimated with very low computational overhead, by leveraging the linearity of the PCE regression model to compute LOO residuals and LOO predictions via analytical, closed-form expressions. The latter are used by the jackknife and jackknife+ CP methods to estimate the predictive interval.
The conformalized PCE method was applied to four different benchmark models, with one, six, seven, and ten input dimensions, respectively. Total-degree PCEs were employed for the multivariate models. Different PCE configurations were used, by varying the maximum polynomial degree, experimental design size, CP method (jackknife versus jackknife+), and non-conformity score type (normalized versus non-normalized). Based on the numerical results, the conformalized PCE is able to provide empirical coverages close to the target level. Training data availability, equivalently, experimental design size is important for reducing the variability in predictive interval width and coverage and stabilizing the PCE. The same can be observed when increasing the PCE maximum polynomial degree, thus increasing its accuracy as well (assuming again sufficient training data). The jackknife method typically results in more conservative conformalized PCEs, with wider predictive intervals and empirical coverages above the target level. Contrarily, the jackknife+ method yields coverages closer to the target level. Non-conformity score normalization does not seem to have a significant impact on predictive interval width and coverage, at least for the examined benchmark models.
The conformalized PCE method developed in this work shows definite promise, however, it can be improved in several aspects. First, this work has considered total-degree PCEs only, thus limiting the selected test-cases to low dimensions and/or polynomial degrees. The integration of CP within sparse and adaptive PCEs would allow to extend the method to more challenging problem settings. Second, in its current form, the conformalized PCE method is applicable to scalar outputs only. In principle, it can be applied elementwise to address multidimensional outputs. However, this would be a suboptimal approach, as it does not account for correlations among the output’s elements and would incur high computational costs with increasing output dimension. The development of a conformalized PCE method suitable for multivariate outputs seems to be a natural follow-up, for example based on the combination of multi-output PCEs [51, 52, 53] and multi-output CP [54, 55]. We aim to address these limitations in dedicated follow-up works.
Author contributions
Dimitrios Loukrezis: Conceptualization, Methodology, Software, Validation, Data curation, Writing - original draft, Writing - review & editing.
Dimitris G. Giovanis: Validation, Data curation, Writing - original draft, Writing - review & editing.
[…]
Appendix A Leave-one-out residuals in closed-form
Let $\widehat{\mu}\left(\mathbf{x}\right)$ denote a PCE regression model as in section 2.1. Removing the $m$-th sample from the training data, we get the LOO design matrix $\mathbf{D}_{\sim m}\in\mathbb{R}^{(M-1)\times K}$ and the LOO output vector $\mathbf{y}_{\sim m}\in\mathbb{R}^{M-1}$. The omitted $m$-th row of the original design matrix $\mathbf{D}$ is denoted with $\mathbf{d}_{m}^{\top}\in\mathbb{R}^{1\times K}$, where $\mathbf{d}_{m}^{\top}=\left(\Psi_{1}\left(\mathbf{x}^{(m)}\right),\dots,\Psi_{K}\left(\mathbf{x}^{(m)}\right)\right)$. Then, the LOO equivalent to the least-squares solution (5) is
$\mathbf{c}_{\sim m}=\left(\mathbf{D}_{\sim m}^{\top}\mathbf{D}_{\sim m}\right)^{-1}\mathbf{D}_{\sim m}^{\top}\mathbf{y}_{\sim m},$ (24)
where it also holds that
$\mathbf{D}_{\sim m}^{\top}\mathbf{D}_{\sim m}$ $=\mathbf{D}^{\top}\mathbf{D}-\mathbf{d}_{m}\mathbf{d}_{m}^{\top},$ (25)
$\mathbf{D}_{\sim m}^{\top}\mathbf{y}_{\sim m}$ $=\mathbf{D}^{\top}\mathbf{y}-\mathbf{d}_{m}y^{(m)}.$ (26)
Let $\mathbf{A}=\left(\mathbf{D}^{\top}\mathbf{D}\right)^{-1}\in\mathbb{R}^{K\times K}$. Using the Sherman-Morrison formula we get
$\left(\mathbf{D}^{\top}\mathbf{D}-\mathbf{d}_{m}\mathbf{d}_{m}^{\top}\right)^{-1}=\mathbf{A}+\frac{\mathbf{A}\mathbf{d}_{m}\mathbf{d}_{m}^{\top}\mathbf{A}}{1-\mathbf{d}_{m}^{\top}\mathbf{A}\mathbf{d}_{m}},$ (27)
assuming that $1-\mathbf{d}_{m}^{\top}\mathbf{A}\mathbf{d}_{m}\neq 0$. Note that $\mathbf{d}_{m}^{\top}\mathbf{A}\mathbf{d}_{m}=h_{mm}$, where $h_{mm}$ are the diagonal elements of the hat matrix $\mathbf{H}\in\mathbb{R}^{M\times M}$ defined in (13). Using expressions (25), (26), and (27), expression (24) can now be written as
$\mathbf{c}_{\sim m}$ $=\left(\mathbf{A}+\frac{\mathbf{A}\mathbf{d}_{m}\mathbf{d}_{m}^{\top}\mathbf{A}}{1-h_{mm}}\right)\left(\mathbf{D}^{\top}\mathbf{y}-\mathbf{d}_{m}y^{(m)}\right)$
$=\mathbf{A}\mathbf{D}^{\top}\mathbf{y}-\mathbf{A}\mathbf{d}_{m}y^{(m)}+\frac{\mathbf{A}\mathbf{d}_{m}\mathbf{d}_{m}^{\top}\mathbf{A}\mathbf{D}^{\top}\mathbf{y}}{1-h_{mm}}-\frac{\mathbf{A}\mathbf{d}_{m}\mathbf{d}_{m}^{\top}\mathbf{A}\mathbf{d}_{m}y^{(m)}}{1-h_{mm}}$
$=\mathbf{c}-\mathbf{A}\mathbf{d}_{m}y^{(m)}+\frac{\mathbf{A}\mathbf{d}_{m}\mathbf{d}_{m}^{\top}\mathbf{c}}{1-h_{mm}}-\frac{\mathbf{A}\mathbf{d}_{m}h_{mm}y^{(m)}}{1-h_{mm}}$
$=\mathbf{c}+\mathbf{A}\mathbf{d}_{m}\left(\frac{\mathbf{d}_{m}^{\top}\mathbf{c}-h_{mm}y^{(m)}}{1-h_{mm}}-y^{(m)}\right)$
$=\mathbf{c}+\mathbf{A}\mathbf{d}_{m}\frac{\mathbf{d}_{m}^{\top}\mathbf{c}-y^{(m)}}{1-h_{mm}}$
$=\mathbf{c}+\mathbf{A}\mathbf{d}_{m}\frac{\widehat{y}^{(m)}-y^{(m)}}{1-h_{mm}}$
$=\mathbf{c}-\mathbf{A}\mathbf{d}_{m}\frac{y^{(m)}-\widehat{y}^{(m)}}{1-h_{mm}},$ (28)
where we have used that $\mathbf{c}=\mathbf{A}\mathbf{D}^{\top}\mathbf{y}$ and $\widehat{y}^{(m)}=\widehat{\mu}\left(\mathbf{x}^{(m)}\right)=\mathbf{d}_{m}^{\top}\mathbf{c}$. The LOO model evaluation at $\mathbf{x}^{(m)}$ is then given by
$\widehat{\mu}_{\sim m}\left(\mathbf{x}^{(m)}\right)=\mathbf{d}_{m}^{\top}\mathbf{c}_{\sim m}=\mathbf{d}_{m}^{\top}\mathbf{c}-\mathbf{d}_{m}^{\top}\mathbf{A}\mathbf{d}_{m}\frac{y^{(m)}-\widehat{y}^{(m)}}{1-h_{mm}}=\widehat{y}^{(m)}-h_{mm}\frac{y^{(m)}-\widehat{y}^{(m)}}{1-h_{mm}}.$ (29)
Therefore, the LOO residual defined in (10) is given by
$r_{m}^{\text{LOO}}$ $=y^{(m)}-\widehat{\mu}_{\sim m}\left(\mathbf{x}^{(m)}\right)$
$=y^{(m)}-\widehat{y}^{(m)}+h_{mm}\frac{\widehat{y}^{(m)}-y^{(m)}}{1-h_{mm}}$
$=\frac{y^{(m)}-\widehat{y}^{(m)}}{1-h_{mm}}$
$=\frac{y^{(m)}-\widehat{\mu}\left(\mathbf{x}^{(m)}\right)}{1-h_{mm}}.$ (30)
Appendix B Leave-one-out predictions in closed-form
Using similar notation as in appendix A, the LOO prediction at a test input $\mathbf{x}^{*}$ is given by
$\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)=\mathbf{d}_{*}^{\top}\mathbf{c}_{\sim m},$ (31)
where $\mathbf{d}_{*}^{\top}=\left(\Psi_{1}\left(\mathbf{x}^{*}\right),\dots,\Psi_{K}\left(\mathbf{x}^{*}\right)\right)$. Substituting the LOO PCE coefficients from (24) into (31), we get
$\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)$ $=\mathbf{d}_{*}^{\top}\left(\mathbf{c}-\mathbf{A}\mathbf{d}_{m}\frac{y^{(m)}-\widehat{y}^{(m)}}{1-h_{mm}}\right)$
$=\mathbf{d}_{*}^{\top}\mathbf{c}-\mathbf{d}_{*}^{\top}\mathbf{A}\mathbf{d}_{m}\frac{y^{(m)}-\widehat{y}^{(m)}}{1-h_{mm}}$
$=\widehat{\mu}\left(\mathbf{x}^{*}\right)-\mathbf{d}_{*}^{\top}\mathbf{A}\mathbf{d}_{m}\frac{y^{(m)}-\widehat{y}^{(m)}}{1-h_{mm}}$
$=\widehat{\mu}\left(\mathbf{x}^{*}\right)-\mathbf{d}_{*}^{\top}\mathbf{A}\mathbf{d}_{m}\frac{y^{(m)}-\widehat{\mu}\left(\mathbf{x}^{(m)}\right)}{1-h_{mm}}$
$=\widehat{\mu}\left(\mathbf{x}^{*}\right)-\mathbf{d}_{*}^{\top}\left(\mathbf{D}^{\top}\mathbf{D}\right)^{-1}\mathbf{d}_{m}\frac{y^{(m)}-\widehat{\mu}\left(\mathbf{x}^{(m)}\right)}{1-h_{mm}}$
$=\widehat{\mu}\left(\mathbf{x}^{*}\right)-\mathbf{d}_{*}^{\top}\left(\mathbf{D}^{\top}\mathbf{D}\right)^{-1}\mathbf{d}_{m}r_{m}^{\text{LOO}},$ (32)
where we used that $\mathbf{A}=\left(\mathbf{D}^{\top}\mathbf{D}\right)^{-1}$, $\widehat{\mu}\left(\mathbf{x}^{*}\right)=\mathbf{d}_{*}^{\top}\mathbf{c}$, and the analytical expression (A) for the LOO residual.